OpenStudy (anonymous):

What is the slope of a line that is 2x-8y-3=0?

OpenStudy (anonymous):

2x-8y-3=0 +3 +3 = 2x-8y=3 -2x -2x = -8y=-2x+3 Divide everything by -8 = y=1/4x - 3/8 ---> y= mx + b ---> m is the slope Slope is 1/4

OpenStudy (anonymous):

[m=1/4\], just like @ivy_green